# Algorithmic Behaviors Across Regions: A Geolocation Audit of YouTube Search for COVID-19 Misinformation between the United States and South Africa

### v1

@@ 1 Introduction
Since March 2020, the World Health Organization (WHO) has designated COVID-19 as a global pandemic. The pandemic continues to pose a public health threat, registering 206K COVID-19 cases and 3.5K deaths worldwide in July 2024 alone (WHO 2024). As the largest video search engine, YouTube has emerged as a vital tool for finding health-related information online, particularly during outbreaks and global pandemics (Dubey et al. 2014; Bora et al. 2018; Khatri et al. 2020). However, YouTube has faced criticism for disseminating COVID-19 misinformation globally to its users, with fact-checkers calling the platform a “major conduit of fake news” (Milmo 2022). This misinformation has undermined public health efforts worldwide, fueling vaccine hesitancy and eroding trust in health institutions (WHO 2022). Thus, there is a pressing need for empirical investigation into search engine systems to ensure algorithmic accountability, safeguard global health, and promote a more responsible, trustworthy web.
@@
We find multiple instances where bots in SA encountered statistically significantly more misinformative SERPs than bots in the US. These disparities were observed within the top-10 search results (p$<$0.001, r=0.49) and search results sorted by YouTube’s “Relevance” filter (p$<$0.001, r=0.86) with medium to large effect sizes, indicating practical significance. Since 95% of user traffic is directed towards the first page of the search results (Kaye 2013) and YouTube employs the “Relevance” filter by default, users in SA may likely encounter significantly more misinformative SERPs than users in the US. Given YouTube’s established importance for finding health information, this could raise personal health risks for users in SA by potentially negatively influencing their beliefs, health practices, and decisions. Overall, our work highlights the contrasting algorithmic behaviors of YouTube’s search function in two countries within the context of COVID-19 misinformation, underscoring the need for YouTube to regulate its algorithmic behaviors consistently across different regions of the Globe.
@@
To our knowledge, our study is the first large-scale geolocation-based comparative audit of YouTube search for COVID-19 misinformation across countries in the Global North and the Global South. This research explores the less-explored context of the Global South, extending the scope of previous algorithmic audits, which primarily centered on the Global North. Our work provides an elaborate understanding of YouTube’s search engine across 4 distinct search filters for 8 misinformation topics. Through a 10-day data collection and extensive labeling process, we amassed a labeled dataset of 915K search results (10,139 unique videos) and trained a classifier to detect COVID-19 misinformation in YouTube videos [^1] We make our dataset and classifier public upon acceptance.. Overall, our work highlights the contrasting algorithmic behaviors of YouTube’s search function in two countries within the context of COVID-19 misinformation, underscoring the need for YouTube to regulate its algorithmic behaviors consistently across different regions of the Globe.
@@ 2 Related Work
Search engines determine what information is relevant and shape user behavior, impacting aspects such as political voting behavior (Diakopoulos et al. 2018), scientific knowledge (Papadamou et al. 2022), and beliefs (Knobloch-Westerwick et al. 2015). Despite their societal importance, search engines operate without external regulation, leaving the credibility of the content unverified. Consequently, researchers have investigated search systems by conducting algorithmic audits that empirically measure and understand the conditions in which problematic content arises on the platform. Several studies have investigated search engines for misinformation (Hussein et al. 2020; Papadamou et al. 2022; Tomlein et al. 2021), conspiracy theories (Chen et al. 2023; Faddoul et al. 2020), hate speech and extremism (Albadi et al. 2022; Ribeiro et al. 2020), and partisanship (Robertson et al. 2018). Among the various methods to audit search engines, we utilize sock-puppet audits (programming bots to emulate real users), a methodology commonly employed in prior audits (Hussein et al. 2020; Papadamou et al. 2022; Kliman-Silver et al. 2015) for its control over experimental variables. Our study adds to the existing sock-puppet audit studies, investigating the prevalence of COVID-19 misinformation on YouTube between the US and SA.
@@
Prior works have highlighted the growing need to consider the Global South in algorithmic audit research (Costanza-Chock et al. 2022). In a recent paper, Urman et al. (2024) argued that most algorithmic audit research is skewed towards the Western context, noting that “countries located outside of North America and Western Europe are understudied.” A limited number of studies have conducted algorithmic audits within the Global South context, examining factors such as language (Narain et al. 2023) and culture (Dammu et al. 2024). Among these audits conducted in the Global South contexts, only a few considered geolocation, with a vast majority focusing on Google Search (Le et al. 2022; Dabran-Zivan et al. 2023). We contribute to the less-researched context of the Global South and extend prior algorithmic audits on YouTube, the largest video search engine.
@@
Given the importance of search engines for finding health-related information during the COVID-19 pandemic (Khatri et al. 2020), several scholars have audited search engines for COVID-19 misinformation (Li et al. 2020; Kravets et al. 2023; Papadamou et al. 2022) and anti-vaccine content (Juneja et al. 2021; Papadamou et al. 2022). In addition, researchers have engineered various features and built machine learning models to automatically detect COVID-19 misinformation online (Ng et al. 2023; Papadamou et al. 2022; Serrano et al. 2020). However, most prior studies have primarily focused on the Global North, leaving a gap in systematic audits of search engines for COVID-19 misinformation in the Global South contexts.
Existing literature regarding the Global South has several gaps, often probing search engines with the default search filter for a single day and focusing on a narrow range of topics. For example, researchers in (Narain et al. 2023) focused on the language aspect of the audit, examining COVID-19 videos in 11 widely spoken languages across Africa on YouTube and collecting 562 videos for evaluation. Dabran-Zivan et al. (2023) examined geolocation and language, analyzing 3 COVID-19 conspiracy topics on Google Search across 4 languages in 10 countries for a single day, collecting 330 search results. We systematically compare the prevalence of COVID-19 misinformation on YouTube across countries in the Global North and Global South. Across geolocations in the US and SA, we examine 8 different COVID-19 misinformation topics and 4 distinct search filters over 10 days of data collection, yielding 915K search results. Unlike previous audits with single geolocations per country, we selected three per country for fine-grained comparative analysis within countries.
@@ 3 Audit Experiment Setup
We considered the US and SA because they were heavily affected by the pandemic in the Global North and the Global South, respectively, making them vulnerable to COVID-19 misinformation [^2] As of May 7th, 2024, the US and SA continue to have the highest reported cases of COVID-19 in the continents of North America and Africa, respectively (WHO 2024). (WHO 2024). Previous work established that Google personalizes search results across different states in the US (Kliman-Silver et al. 2015). Therefore, we chose to identify three states in each country to achieve a more robust analysis of COVID-19 misinformation in YouTube search results of the selected countries. We identified three states in the US and three provinces [^3] South Africa operates on a provincial government system, unlike the state government system in the US. in South Africa with the highest total confirmed cases of COVID-19 (Adeline et al. 2020; Wits University 2023), which may make them susceptible to COVID-19 misinformation. To capture the highest proportion of the population in these states and provinces, we selected the largest populated city as the geolocation. For US, we selected Los Angeles (California), Houston (Texas), and Jacksonville (Florida). For South Africa, we selected Johannesburg (Gauteng), Durban (KwaZulu-Natal), and Cape Town (Western Cape). Figure 1 depicts the geolocations chosen in the US and SA.
@@
To curate globally persistent COVID-19 misinformation topics, we partnered with expert fact-checkers at Pesacheck [^4] https://www.pesacheck.org. As Africa’s largest indigenous fact-checking organization, PesaCheck is affiliated with the International Fact-Checking Network (ICFN) and collaborates with expert fact-checkers worldwide. Since our audit study investigates both SA and the US, our partnership enabled us to incorporate their expertise on the COVID-19 misinformation widely circulating globally into our study. Utilizing editorial coverages and frequent fact-checking, the expert fact-checkers identified 8 globally persistent COVID-19 misinformation topics, providing a dataset of 362 YouTube videos related to these topics. Please refer to Appendix Table 4 for the sample YouTube videos provided by PesaCheck. These videos originate from channels associated with 29 countries spanning the Global North and the Global South (See Appendix Table 3). Table 1 presents the 8 topics and samples of our curated search queries, which we explain next.
@@
We utilized three methods to curate search queries for each topic. Given that English is the only language spoken commonly in both the US and SA [^5] English is one of eleven official languages in SA, with 10% of people speaking English natively at home and is most commonly spoken in business and commerce (Doochin 2019)., we focused on queries in English, thus enabling a controlled cross-country comparison in the amount of COVID-19 misinformation returned by YouTube’s search algorithm.
@@
First, we utilized YouTube video tags in the videos provided by Pesacheck. Video tags are descriptive keywords representing how content creators want their videos to be discovered (Google 2022). Misinformative videos often contain tags describing misleading narratives, effectively serving as potential search queries for identifying more misinformative videos (Juneja et al. 2023). We collected 2,911 video tags from the videos and applied systematic filtering criteria employed in prior work (Juneja et al. 2021). We excluded queries mentioning individuals or news organizations (e.g., “Obama”, “Republic TV”), overly generic terms (e.g., “breaking news”), excessively specific terms (e.g., “COVID3rdWaveInMyanmar”), and irrelevant terms (e.g., “cute puppies”). Additionally, we manually merged similar queries (e.g. “man-made” and “man-made virus” into the single query “man-made virus”). These systematic filtering steps resulted in 48 tags. For example, the tags included problematic terms regarding virus origination in China (“Kungflu,” “China virus”) and terms on self-treatment methods (“local concoction”).
@@
Second, for each topic, we identified important keywords found in video titles, descriptions, and transcripts from PesaCheck-provided videos. We created a document for each topic by concatenating the video metadata and applying standard preprocessing steps. To extract keywords from each topic document, we used KeyBERT (Grootendorst 2022), a keyword extraction tool that requires two inputs: an embedding model and a document. First, KeyBERT applies the embedding model on the topic document to generate the document embeddings and extract word embeddings for N-gram phrases [^6] We limited our phrases to unigram, bigram, and trigrams.. Then, KeyBERT uses cosine similarity to identify keywords with the highest similarity with the document itself, providing words that best describe the entire document. We chose the Sentence Transformer all-mpnet-base-v2 due to their best performance in generating sentence embeddings (Song et al. 2020). Since the model was not trained on COVID-19-related texts, we fine-tuned the model for domain adaptation using COVID-19 question-answering datasets (Möller et al. 2020; Lu 2022; Tang et al. 2020). See Appendix A for fine-tuning details. After fine-tuning, we used KeyBERT on each of the 8 topic documents to extract 5 keywords per document, resulting in 40 search queries.
@@
Third, we included the 49 search keywords provided by PesaCheck because the keywords were curated by expert fact-checkers to find misinformative videos. Combining the search queries from all three methods, we obtained 137 search queries. Then, we manually removed duplicate queries, combined similar queries, and randomly selected 6 search queries per topic, resulting in a final set of 48 search queries (see Appendix Table 5). To ensure that the search queries output search results relevant to the COVID-19 pandemic, we identified keyword variations of COVID-19 and employed YouTube search operators, which are special commands that can be utilized to efficiently refine their searches on YouTube (Zarr 2020). We selected 8 keyword variations of COVID-19 using a COVID-19 Twitter dataset (Chen et al. 2020). Each query was formatted as:
@@
To control for possible confounding factors that may affect our audit, we followed standard noise control procedures based on prior work (Kliman-Silver et al. 2015; Hussein et al. 2020; Juneja et al. 2021). To differentiate between noise and geolocation-based personalization in SERPs, we created identical twin bots, consisting of a treatment bot and its corresponding control bot that conducted the same actions simultaneously at each geolocation. Thus, any difference in the search results between the twin bots should be attributed to noise rather than personalization [^7] Note that YouTube’s search engine is a black box. Even after controlling for all known sources of noise, there could be some sources of noise we are unaware of.. If the differences in the SERPs between two geolocations exceeded the noise, it can only be attributed to geolocation-based personalization. In our experiment, we placed twin bots at each geolocation, resulting in 12 bots.
@@
Changing IP addresses using VMs (Hussein et al. 2020) and proxies (Boeker et al. 2022) are common methods to conduct geolocation audits, especially in the US context. However, obtaining fine-grained coverage of IP addresses through VMs and proxy services in South Africa was challenging. The only AWS coverage in South Africa, let alone in Africa, is in Cape Town. Meanwhile, nearly all the proxy and VPN providers except IPRoyal were unfeasible due to the steep pricing or the lack of proxy coverage of our selected geolocations in South Africa (see Appendix Table 6 for further details). Due to challenges associated with IP addresses, we initially turned to the “geospoofing” method used in Kliman-Silver et al. (2015), in which they fed precise latitude and longitude coordinate information to automated scripts and obtained personalized search results, providing a cost-effective alternative. For brevity, we summarized the results from our validation experiments and left the details in the Appendix. We defined the metric for geolocation-based personalization in Appendix B. In our first validation experiment, we curated a set of search queries and validated that the queries elicited geolocation-based personalization, resulting in personalized SERPs based on geolocation (Appendix C). Using the validated search queries, we performed a second validation experiment and found that YouTube uses IP geolocation instead of geospoofed location to personalize search results (Appendix D). Thus, we conducted a third validation experiment using IPRoyal proxies, which validated the accuracy and consistency of the proxies in giving us the correct IP geolocations for our experiment (Appendix E).

### v2

@@ 1 Introduction
Since March 2020, the World Health Organization (WHO) has designated COVID-19 as a global pandemic. The pandemic continues to pose a public health threat, registering 206K COVID-19 cases and 3.5K deaths worldwide in July 2024 alone (WHO 2024). As the largest video search engine, YouTube has emerged as a vital tool for finding health-related information online, particularly during outbreaks and global pandemics (Dubey et al. 2014; Bora et al. 2018; Khatri et al. 2020). However, YouTube has faced criticism for disseminating COVID-19 misinformation globally to its users, with fact-checkers calling the platform a “major conduit of fake news” (Milmo 2022). This misinformation has undermined public health efforts worldwide, fueling vaccine hesitancy and eroding trust in health institutions (Do Nascimento et al. 2022). Thus, there is a pressing need for empirical investigation of search engine systems to ensure algorithmic accountability, safeguard global health, and promote a more responsible and trustworthy web.
@@
We find multiple instances where bots in SA encountered statistically significantly more misinformative SERPs than bots in the US. These disparities were observed within the top-10 search results (p$<$0.001, r=0.49) and search results sorted by YouTube’s “Relevance” filter (p$<$0.001, r=0.86) with medium to large effect sizes, indicating practical significance. Since 95% of the user traffic is directed towards the first page of the search results (Kaye 2013) and YouTube employs the “Relevance” filter by default, users in SA may likely encounter significantly more misinformative SERPs than users in the US. Given YouTube’s established importance for finding health information, this could raise personal health risks for users in SA by potentially negatively influencing their beliefs, health practices, and decisions. Overall, our work highlights the contrasting algorithmic behaviors of YouTube’s search function in two countries within the context of COVID-19 misinformation, underscoring the need for YouTube to regulate its algorithmic behaviors consistently across different regions of the Globe.
@@
To our knowledge, our study is the first large-scale geolocation-based comparative audit of YouTube search for COVID-19 misinformation across countries in the Global North and the Global South. This research explores the less-explored context of the Global South, extending the scope of previous algorithmic audits, which primarily centered on the Global North. Our work provides an elaborate understanding of YouTube’s search engine across 4 distinct search filters for 8 misinformation topics. Through a 10-day data collection and extensive labeling process, we amassed a labeled dataset [^1] https://github.com/social-comp/YouTubeAuditGeolocation-data of 915K search results (10,139 unique videos) and trained a classifier [^2] https://huggingface.co/SocialCompUW/youtube-covid-misinfo-detect to detect COVID-19 misinformation in YouTube videos. Overall, our work highlights the contrasting algorithmic behaviors of YouTube’s search function in two countries within the context of COVID-19 misinformation, underscoring the need for YouTube to regulate its algorithmic behaviors consistently across different regions of the Globe.
@@ 2 Related Work
Search engines determine what information is relevant and shape user behavior, impacting aspects such as political voting behavior (Diakopoulos et al. 2018), scientific knowledge (Papadamou et al. 2022), and beliefs (Knobloch-Westerwick et al. 2015). Despite their societal importance, search engines operate without external regulation, leaving the credibility of content unverified. Consequently, researchers have investigated search systems by conducting algorithmic audits that empirically measure and understand the conditions under which problematic content arises on the platform. Several studies have investigated search engines for misinformation (Hussein, Juneja, and Mitra 2020; Papadamou et al. 2022; Tomlein et al. 2021), conspiracy theories (Chen et al. 2023; Faddoul et al. 2020), hate speech and extremism (Albadi, Kurdi, and Mishra 2022; Ribeiro et al. 2020), and partisanship (Robertson et al. 2018). Among the various methods for auditing search engines, we utilize sock-puppet audits (programming bots to emulate real users), a methodology commonly employed in prior audits (Hussein, Juneja, and Mitra 2020; Papadamou et al. 2022; Kliman-Silver et al. 2015) for its control over experimental variables. Our study adds to the existing sock-puppet audit studies, investigating the prevalence of COVID-19 misinformation on YouTube between the US and SA.
@@
Prior works have highlighted the growing need to consider the Global South in algorithmic audit research (Costanza-Chock, Raji, and Buolamwini 2022). In a recent paper, Urman, Makhortykh, and Hannák (2024) argued that most algorithmic audit research is skewed towards the Western context, noting that “countries located outside of North America and Western Europe are understudied.” A limited number of studies have conducted algorithmic audits within the Global South context, examining factors such as language (Narain et al. 2023) and culture (Dammu et al. 2024). Among these audits conducted in the Global South contexts, only a few considered geolocation, with a vast majority focusing on Google Search (Le et al. 2022; Dabran-Zivan et al. 2023). We contribute to the less-researched context of the Global South and extend prior algorithmic audits on YouTube, the largest video search engine.
@@
Given the importance of search engines for finding health-related information during the COVID-19 pandemic (Khatri et al. 2020), several scholars have audited search engines for COVID-19 misinformation (Li et al. 2020; Kravets et al. 2023; Papadamou et al. 2022) and anti-vaccine content (Juneja and Mitra 2021; Papadamou et al. 2022). In addition, researchers have engineered various features and built machine learning models to automatically detect COVID-19 misinformation online (Ng, Hoffmann Pham, and Luengo-Oroz 2023; Papadamou et al. 2022; Medina Serrano, Papakyriakopoulos, and Hegelich 2020). However, most prior studies have primarily focused on the Global North, leaving a gap in audits of search engines for COVID-19 misinformation in the Global South contexts.
Existing literature regarding the Global South has several gaps, often probing search engines with the default search filter for a single day and focusing on a narrow range of topics. For example, researchers in Narain et al. (2023) focused on the language aspect of the audit, examining COVID-19 videos in 11 widely spoken languages across Africa on YouTube and collecting 562 videos for evaluation. Dabran-Zivan et al. (2023) examined geolocation and language, analyzing 3 COVID-19 conspiracy topics on Google Search across 4 languages in 10 countries for a single day, collecting 330 search results. We systematically compare the prevalence of COVID-19 misinformation on YouTube across countries in the Global North and Global South. Across geolocations in the US and SA, we examine 8 different COVID-19 misinformation topics and 4 distinct search filters over 10 days of data collection, yielding 915K search results. Unlike previous audits with single geolocations per country, we selected three per country for fine-grained comparative analysis within countries.
@@ 3 Audit Experiment Setup
We considered the US and SA because they were heavily affected by the pandemic in the Global North and the Global South, respectively, making them vulnerable to COVID-19 misinformation [^3] As of May 7th, 2024, the US and SA continue to have the highest reported cases of COVID-19 in the continents of North America and Africa, respectively (WHO 2024). (WHO 2024). Previous work established that Google personalizes search results across different states in the US (Kliman-Silver et al. 2015). Therefore, we chose to identify three states in each country to achieve a more robust analysis of COVID-19 misinformation in YouTube search results of the selected countries. We identified three states in the US and three provinces [^4] South Africa operates on a provincial government system. in South Africa with the highest total confirmed cases of COVID-19 (Adeline et al. 2020; Wits University 2023), which may make them susceptible to COVID-19 misinformation. To capture the highest proportion of the population in these states and provinces, we selected the largest populated city as the geolocation. For US, we selected Los Angeles (California), Houston (Texas), and Jacksonville (Florida). For South Africa, we selected Johannesburg (Gauteng), Durban (KwaZulu-Natal), and Cape Town (Western Cape). Figure 1 depicts the geolocations chosen in the US and SA.
@@
To curate globally persistent COVID-19 misinformation topics, we partnered with expert fact-checkers at Pesacheck. [^5] https://www.pesacheck.org As Africa’s largest indigenous fact-checking organization, PesaCheck is affiliated with the International Fact-Checking Network (IFCN) and collaborates with expert fact checkers worldwide, making them well equipped to identify misinformation that circulates in diverse regions worldwide.
Using editorial coverage, frequently fact-checked information and social media analytics tools, the expert fact-checkers identified eight globally persistent COVID-19 misinformation topics and provided a dataset of 362 fact-checked YouTube videos related to these topics. These videos originate from channels associated with 29 countries, spanning the Global North and the Global South, including Algeria, China, South Africa, and the United States. This diverse representation–18 countries from the Global North and 11 from the Global South (see Appendix Table 3)–ensures that the search queries developed from these videos are representative across different geographic regions. Refer to Appendix Table 4 for sample YouTube videos provided by PesaCheck. Table 1 presents the 8 topics and samples of our curated search queries, which we explain next.
@@
We utilized three methods to curate search queries for each topic. Given that English is the only language spoken commonly in both the US and SA, [^6] English is one of eleven official languages in SA, with 10% of people speaking English natively at home and is most commonly spoken in business and commerce (Doochin 2019). we focused on queries in English, allowing for a controlled cross-country comparison in the amount of COVID-19 misinformation returned by YouTube’s search algorithm.
@@
First, we used YouTube video tags in the videos provided by Pesacheck. Video tags are descriptive keywords that represent how content creators want their videos to be discovered (Google 2022). Misinformative videos often contain tags that describe misleading narratives, effectively serving as potential search queries to identify more misinformative videos (Juneja, Bhuiyan, and Mitra 2023). We collected 2,911 video tags from the videos and applied systematic filtering criteria used in previous work (Juneja and Mitra 2021). We excluded queries mentioning individuals or news organizations (e.g., “Obama”, “Republic TV”), overly generic terms (e.g., “breaking news”), excessively specific terms (e.g., “COVID3rdWaveInMyanmar”), and irrelevant terms (e.g., “cute puppies”). Additionally, we manually merged similar queries (e.g. “man-made” and “man-made virus” into the single query “man-made virus”). These systematic filtering steps resulted in 48 tags. For example, the tags included problematic terms regarding the origin of the virus in China (“Kungflu,” “China virus”) and terms on self-treatment methods (“local concoction”).
@@
Second, for each topic, we identified important keywords found in video titles, descriptions, and transcripts from PesaCheck-provided videos. We created a document for each topic by concatenating the video metadata and applying standard preprocessing steps. To extract keywords from each topic document, we used KeyBERT (Grootendorst 2022), a keyword extraction tool that requires two inputs: an embedding model and a document. First, KeyBERT applies the embedding model to the topic document to generate the document embeddings and extract word embeddings for N-gram phrases. [^7] We limited our phrases to unigram, bigram, and trigrams. Then, KeyBERT uses cosine similarity to identify keywords with the highest similarity with the document itself, providing words that best describe the entire document. We chose the Sentence Transformer all-mpnet-base-v2 due to their best performance in generating sentence embeddings (Song et al. 2020). Since the model was not trained on COVID-19-related texts, we fine-tuned the model for domain adaptation using COVID-19 question-answering datasets (Möller et al. 2020; Lu 2022; Tang et al. 2020). See Appendix A for fine-tuning details. After fine-tuning, we used KeyBERT on each of the 8 topic documents to extract 5 keywords per document, resulting in 40 search queries.
@@
Third, we included the 49 search keywords provided by PesaCheck because the keywords were curated by expert fact-checkers to find misinformative videos. Combining the search queries from all three methods, we obtained 137 search queries. Then, we manually removed duplicate queries, combined similar queries, and randomly selected 6 search queries per topic, resulting in a final set of 48 search queries (see Appendix Table 5). To ensure that the search queries output search results relevant to the COVID-19 pandemic, we identified keyword variations of COVID-19 and employed YouTube search operators, which are special commands that can be utilized to efficiently refine their searches on YouTube (Zarr 2020). We selected 8 keyword variations of COVID-19 using a COVID-19 Twitter dataset (Chen, Lerman, and Ferrara 2020). Each query was formatted as:
@@
To control for possible confounding factors that may affect our audit, we followed standard noise control procedures based on prior work (Kliman-Silver et al. 2015; Hussein, Juneja, and Mitra 2020; Juneja and Mitra 2021). To differentiate between noise and geolocation-based personalization in SERPs, we created identical twin bots, consisting of a treatment bot and its corresponding control bot that conducted the same actions simultaneously at each geolocation. Thus, any difference in the search results between the twin bots should be attributed to noise rather than personalization. [^8] Note that YouTube’s search engine is a black box. Even after controlling for all known sources of noise, there could be some sources of noise we are unaware of. If the differences in the SERPs between two geolocations exceeded the noise, it can only be attributed to geolocation-based personalization. In our experiment, we placed twin bots at each geolocation, resulting in 12 bots.
@@
Changing IP addresses using VMs (Hussein, Juneja, and Mitra 2020) and proxies (Boeker and Urman 2022) are common methods to conduct geolocation audits, especially in the US context. However, obtaining fine-grained coverage of IP addresses through VMs and proxy services in South Africa was challenging. The only AWS coverage in South Africa, let alone in Africa, is in Cape Town. Meanwhile, nearly all the proxy and VPN providers except IPRoyal were unfeasible due to the steep pricing or the lack of proxy coverage of our selected geolocations in South Africa (see Appendix Table 6 for further details). Due to challenges associated with IP addresses, we initially turned to the “geospoofing” method used in Kliman-Silver et al. (2015), in which they fed precise latitude and longitude coordinate information to automated scripts and obtained personalized search results, providing a cost-effective alternative. For brevity, we summarized the results from our validation experiments and left the details in the Appendix. We defined the metric for geolocation-based personalization in Appendix B. In our first validation experiment, we curated a set of search queries and validated that the queries elicited geolocation-based personalization, resulting in personalized SERPs based on geolocation (Appendix C). Using the validated search queries, we performed a second validation experiment and found that YouTube uses IP geolocation instead of geospoofed location to personalize search results (Appendix D). Thus, we conducted a third validation experiment using IPRoyal proxies, which validated the accuracy and consistency of the proxies in giving us the correct IP geolocations for our experiment (Appendix E).